OpenStudy (anonymous):

y = x −(256/x^2) i need to fine the interecepts,realatiev extremas,point sof imflection and asymptote. PleaSE any help would be apreciated

OpenStudy (anonymous):

\[f(x)=x-\frac{256}{x^2}\]

OpenStudy (anonymous):

\[f'(x)=1+\frac{512}{x^3}=\frac{x^3+512}{x^3}\]

OpenStudy (anonymous):

first asymptotes. obviously one at x = 0 because this is undefined there. a slant asymptote at \[y=x\]

OpenStudy (anonymous):

ok now set the derivative = 0 and solve to get critical points. i.e. set \[x^3+512=0\] \[x^3=-512\] \[x=-8\] that is your critical point

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

it is a relative max

OpenStudy (anonymous):

we know this because the derivative is positive to the left of -8 and negative to the right, which means your function was increasing and then decreasing. hence a max

OpenStudy (anonymous):

to find inflection point take derivative again. don't use the quotient rule, start with \[f'(x)=1+\frac{215}{x^3}\] derivative is \[f''(x)=-\frac{1536}{x^4}\]

OpenStudy (anonymous):

this thing is obviously negative for all values of x because the denominator is positive always and you have a minus sign out front, so function is always concave down

OpenStudy (anonymous):

this also shows that -8 is a relative max, because the second derivative is negative there

OpenStudy (anonymous):

and that is all except remember that -8 is not the relative max, it is the x value that gives the relative max. the relative max is what you get when you evaluate the function there. i will let you do that for yourself

OpenStudy (anonymous):

if you want a nice picture of your function look here http://www.wolframalpha.com/input/?i=y%3Dx-256%2Fx^2

OpenStudy (anonymous):

thanks satelite, so what about the interecpt..r they none?and also theres no relative minimum then n only a maximum right?and to fine the maximum i would just plug -8 ninto the second derivitve right

OpenStudy (anonymous):

thanks satelite, so what about the interecpt..r they none?and also theres no relative minimum then n only a maximum right?and to fine the maximum i would just plug -8 ninto the second derivitve right

OpenStudy (anonymous):

no no, for maximum you want maximum of the function!

OpenStudy (anonymous):

plug -8 into the function, not the derivative or the second derivative

OpenStudy (anonymous):

yes there is no relative min. only one critical point

OpenStudy (anonymous):

there is no y intercept since this has an asymptote at x = 0 (the y axis) it never touches it

OpenStudy (anonymous):

the x intercept is where the function is 0. set it = 0 and solve

OpenStudy (anonymous):

it is zero if\[ x = \sqrt[3]{256}\]